OpenStudy (anonymous):

Color of Flowers Number Red 8 Orange 2 Green 5 Yellow 5 A coin Ted selects a flower from the box randomly without looking and then tosses the fair coin. What is the probability that he will select a red flower and that the coin will land tails up? 1 over 10 1 over 5 2 over 5 4 over 5

OpenStudy (anonymous):

i think A

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

how many flowers are there?

OpenStudy (misty1212):

oh nvm there are 20 did you do \[\frac{8}{20}\times \frac{1}{2}\]?

OpenStudy (anonymous):

no you have to simplify the 8 over 20

OpenStudy (anonymous):

to 2/5

OpenStudy (misty1212):

yes to get the final answer sure

OpenStudy (anonymous):

so its A

OpenStudy (misty1212):

\[\frac{8}{20}=\frac{2}{5}\] but \[\frac{2}{5}\times \frac{1}{2}=\frac{1}{5}\] right?

OpenStudy (anonymous):

no 5 x 2 is 10 not 5

OpenStudy (misty1212):

???

OpenStudy (misty1212):

\[\frac{2}{5}\times \frac{1}{2}=\frac{1}{5}\] for sure !